# An iterative method for the solution of Laplace-like equations in high and very high space dimensions

### v3

@@ 3 A moment-based analysis
Let $x\to\|x\|^{2\ell}\,W(x)$ be rotationally symmetric and integrable. Then
$\frac{1}{n\nu_{n}}\int_{S^{n-1}}\chi(\eta)\,\mathrm{d}\eta=\,C(\ell)\int_{\mathbb{R}^{n}}\chi(x)W(x)\,\mathrm{d}x$ (3)
holds for all functions $\chi$ that are positively homogeneous of degree $2\ell\geq 0$, where the constant $C(\ell)$ is given by the expression
$\frac{1}{C(\ell)}=\,n\nu_{n}\int_{0}^{\infty}\!W(re)r^{2\ell+n-1}\,\mathrm{d}r$ (4)
and $e$ is an arbitrarily given unit vector.
Proof
As by assumption $\chi(r\eta)=r^{2\ell}\chi(\eta)$ and $W(r\eta)=W(re)$ holds for $\eta\in S^{n-1}$ and $r>0$, the decomposition (13) leads to
$\int_{\mathbb{R}^{n}}\chi(x)W(x)\,\mathrm{d}x\,=\int_{S^{n-1}}\chi(\eta)\left(\int_{0}^{\infty}\!r^{2\ell}\,W(re)r^{n-1}\,\mathrm{d}r\right)\!\,\mathrm{d}\eta,$
yielding the proposition after some rearrangement. ∎
The choice of the function $W$ can be adapted to the needs. For example, it can be the characteristic function of the unit ball or the normed Gauss function
$W(x)=\left(\frac{1}{\sqrt{\pi}}\right)^{n}\exp\big{(}-\|x\|^{2}\big{)},$ (5)
which breaks down into a product of univariate functions. Another almost trivial, but often very useful observation concerns coordinate changes.
Lemma 4
Under the assumptions from the previous lemma,
$\int_{\mathbb{R}^{n}}\chi(x)W(x)\,\mathrm{d}x=\int_{\mathbb{R}^{n}}\chi(Qx)W(x)\,\mathrm{d}x$ (6)
holds for all orthogonal $(n\times n)$-matrices $Q$.
Proof
The transformation theorem for multivariate integrals leads to
$\int_{\mathbb{R}^{n}}\chi(x)W(x)\,\mathrm{d}x=\int_{\mathbb{R}^{n}}\chi(Qx)W(Qx)|\mathrm{det}Q|\,\mathrm{d}x.$
Because $W(Qx)=W(x)$ and $\mathrm{det}Q=\pm 1$, the proposition follows. ∎
The expected value and the variance of the function (1), the central moments
$E=\frac{1}{n\nu_{n}}\int_{S^{n-1}}\!\|T^{t}\eta\|^{2}\,\mathrm{d}\eta,\quad V=\frac{1}{n\nu_{n}}\int_{S^{n-1}}\!\big{(}\|T^{t}\eta\|^{2}-E\big{)}^{2}\,\mathrm{d}\eta,$ (7)
are of fundamental importance for our considerations.
Lemma 5
The expected value and the variance of the function (1) depend only on the singular values $\sigma_{1},\ldots,\sigma_{m}$ of the matrix $T$. In terms of the power sums
@@
Inserting the matrix $Q$ from the singular value decomposition $T=Q\Sigma\widetilde{Q}^{t}$ of the matrix $T$, Lemma 3 and Lemma 4 lead to the representation
$\frac{1}{n\nu_{n}}\int_{S^{n-1}}\big{(}\|T^{t}\eta\|^{2})^{k}\,\mathrm{d}\eta=C(k)\int_{\mathbb{R}^{n}}\bigg{(}\sum_{i=1}^{m}\!\sigma_{i}^{2}x_{i}^{2}\bigg{)}^{k}W(x)\,\mathrm{d}x$
of the moment of order $k$ as a homogeneous symmetric polynomial in the $\sigma_{i}^{2}$. With the function $W$ from (5), the volume integral on the right-hand side splits into a sum of products of one-dimensional integrals. Using
$\int_{0}^{\infty}t^{j}\mathrm{e}^{-t^{2}}\,\mathrm{d}t=\frac{1}{2}\,\Gamma\left(\frac{j+1}{2}\right),\;\;\nu_{n}=\frac{2}{n}\,\frac{\pi^{n/2}}{\Gamma(n/2)},$
this integral and the constant $C(k)$ can be calculated in principle. For the first two moments, this is easily possible. For moments of order three and higher, one can take advantage of the fact that the symmetric polynomials are polynomials in the power sums of the $\sigma_{i}^{2}$; see VanDerWarden , or Sturmfels for a more recent treatment. ∎
@@
$\frac{n-m}{n+2}\,\frac{2}{m}\leq V<\frac{2n-2}{n+2}$ (12)
and can be arbitrarily close to the value two for large $n$. The spectral norms of the given matrices $T^{t}$, their maximum singular values $\sigma_{i}$, spread over the interval
$\sqrt{\frac{n}{m}}\leq\|T^{t}\|<\sqrt{n}.$ (13)
However, the variances are most likely of the order $\mathcal{O}(1/m)$.
@@
of the variances $V$ tends to zero as the dimension $m$ increases, but also that the variances cluster the more around their expected value the larger $m$ is. This observation is supported by simple, easily reproducible experiments. Uniformly distributed points on the unit sphere $S^{m-1}$ can be generated from vectors in $\mathbb{R}^{m}$ with components that are subject to the standard normal distribution. Such vectors then follow themselves the standard normal distribution in the $m$-dimensional space. Scaling them to length one gives the desired uniformly distributed points on the unit sphere. This allows one to sample the random variable $X$ for any given dimension $m$.
@@
of the matrices $S$ and $S^{2}$, and can therefore be computed without recourse to the singular values of $T$. Consider the $(n\times m)$-matrix $T$ assigned to an arbitrarily given undirected graph with $m$ vertices and $n-m$ edges that maps the components $x_{i}$ of a vector $x\in\mathbb{R}^{m}$ first to themselves and then to the $n-m$ weighted differences [^1] The square root is important. Why, is explained in Sect. 5.
@@
assigned to the edges of the graph connecting the vertices $i$ and $j$. In quantum physics, matrices of the given kind can be associated with the interaction of particles. They formed the original motivation for this work. In the given case, the matrix $S$ has the form $S=I+L/2$, where $I$ is the identity matrix and $L$ is the Laplacian matrix of the graph. The off-diagonal entries of $L$ are $L_{ij}=-1$ if the vertices $i$ and $j$ are connected by an edge and $L_{ij}=0$ otherwise. The diagonal entries $L_{ii}=d_{i}$ are the vertex degrees, the numbers of edges that meet at the vertices $i$.
@@
For a matrix $T$ of dimension $n\times m$ associated with a graph with $m$ vertices and $n-m$ edges, the expected value and the variance (7) are
@@
where the quantity $d^{2}$ is the mean value of the squares of the vertex degrees:
$d^{2}=\frac{1}{m}\sum_{i=1}^{m}d_{i}^{2}.$ (21)
Proof
By (18), the constants (8) possess the representation
$A_{1}=n,\quad A_{2}=\frac{d^{2}-6}{4}\,m\,+\,\frac{5}{2}\,n$
in the case considered here. The proposition therefore follows from (9). ∎
If the squares of the vertex degrees remain bounded in the mean, the variance tends to zero inversely proportional to the total number of vertices and edges. The matrices assigned to graphs whose vertices up to one are connected with a designated central vertex, but not to each other, form the other extreme. For these matrices, the variances decrease to a limit value greater than zero as the number of vertices increases. However, such matrices are a very rare exception, not only with respect to the above random matrices, but also in the context of matrices assigned to graphs. Consider the random graphs with a fixed number $m$ of vertices that are with a given probability $p$ connected by an edge, or those with a fixed number $m$ of vertices and $n-m$ of edges. Sampling the variances assigned to a large number of graphs in such a class, one sees that these variances exceed the value $2/(m+1)$ with a very high probability at most by a factor that is only slightly greater than one, if at all. If all vertices are connected with each other, if their degree is $d_{i}=m-1$ and the number of edges is therefore $n-m=m(m-1)/2$, the variance is
$V=\frac{m^{2}+m-2}{m^{2}+m+4}\frac{2}{m+1}.$ (22)
With the exception of a few extreme cases, it can be observed that the distribution of the values $\|T^{t}\eta\|^{2}$ for points $\eta$ on the unit sphere of $\mathbb{R}^{n}$ more and more approaches the normal distribution with the expected value and the variance (7) as the dimensions increase, a fact that underlines the importance of these quantities. This can be seen by evaluating the expression $\|T^{t}\eta\|^{2}$ at a large number of independently chosen, uniformly distributed points $\eta$ on the unit sphere and comparing the frequency distribution of the resulting values with the given Gauss function
$\frac{1}{\sqrt{2\pi V}}\,\exp\left(-\,\frac{(x-E)^{2}}{2V}\,\right).$ (23)
Let the matrix $T$ have the singular value decomposition $T=Q\Sigma\widetilde{Q}^{t}$. The frequency distribution of the values $\|\Sigma^{t}\eta\|^{2}=\|T^{t}Q\eta\|^{2}$ on the unit sphere then coincides with the distribution of the values $\|T^{t}\eta\|^{2}$. Let $\Sigma_{0}$ be the $(m\times m)$-diagonal matrix that is composed of the singular values. Given the above remarks about the generation of uniformly distributed points on a unit sphere, sampling the values $\|T^{t}\eta\|^{2}$ for a large number of points $\eta\in S^{n-1}$ means to sample the ratio
@@
for a large number of vectors $x\in\mathbb{R}^{m}$ and $y\in\mathbb{R}^{n-m}$ with standard normally distributed components. The values $\|y\|^{2}$ are then distributed according to the $\chi^{2}$-distribution with $n-m$ degrees of freedom. The direct calculation of $\|y\|^{2}$ can thus be replaced by the calculation of a single scalar quantity following this distribution. The amount of work then remains strictly proportional to the dimension $m$, no matter how large the difference $n-m$ of the dimensions is. Let $T$ be the matrix assigned to the graph associated with the $\mathrm{C}_{60}$-fullerene molecule, which consists of the ninety edges of a truncated icosahedron and its sixty corners as vertices. The degree of these vertices is three and the assigned variance therefore $V=9/380$. The frequency distribution of the values $\|T^{t}\eta\|^{2}$ for a million randomly chosen points $\eta$ on the unit sphere and the corresponding Gauss function (23) are shown in Fig. 1.

### v5

@@ 3 The angular distribution
The expected value and the variance (14) depend only on the singular values $\sigma_{1},\ldots,\sigma_{m}$ of the matrix $T$. In terms of the power sums
@@
[…]
$\frac{1}{n\nu_{n}}\int_{S^{n-1}}\big{(}\|T^{t}\eta\|^{2})^{k}\,\mathrm{d}\eta=C(2k)\int_{\mathbb{R}^{n}}\bigg{(}\sum_{i=1}^{m}\!\sigma_{i}^{2}\omega_{i}^{2}\bigg{)}^{k}W(\omega)\,\mathrm{d}\omega$
of the moment of order $k$ as a homogeneous symmetric polynomial of degree $k$ in the $\sigma_{i}^{2}$. The volume integral on the right-hand side splits into a sum of products of one-dimensional integrals. In principle, it can be calculated this way. For the moments of order one and two, this is possible without problems. For higher-order moments, the number of terms to be considered separately increases rapidly. One can then take advantage of the fact that the symmetric polynomials are polynomials in the power sums of the $\sigma_{i}^{2}$; see VanDerWarden , or Sturmfels for a more recent treatment. ∎
@@
$\frac{n-m}{n+2}\,\frac{2}{m}\leq V<\frac{2n-2}{n+2}.$ (19)
However, they are most likely of the order $\mathcal{O}(1/m)$.
@@
of the variances (18) tends to zero as the dimension $m$ increases, but also that the variances increasingly cluster around their expected value as $m$ increases. This observation is supported by simple experiments. Uniformly distributed points on the unit sphere $S^{m-1}$ can be generated from vectors in $\mathbb{R}^{m}$ with independent, standard normally distributed components. Such vectors themselves follow the standard normal distribution in the $m$-dimensional space. Scaling them to length one gives the desired uniformly distributed points on the unit sphere. This allows one to sample the random variable $X$ for any given dimension $m$.
@@
of the matrices $S$ and $S^{2}$, and can therefore be computed without recourse to the singular values of $T$. Consider the $(n\times m)$-matrix $T$ assigned to an arbitrarily given undirected graph with $m$ vertices and $n-m$ edges that maps the components $x_{i}$ of a vector $x\in\mathbb{R}^{m}$ first to themselves and then to the $n-m$ weighted differences [^1] The square root is important. Why, is explained in Sect. 4.
@@
assigned to the edges of the graph connecting the vertices $i$ and $j$. In quantum physics, matrices of the given kind can be associated with the interaction of particles. In the given case, the matrix $S$ has the form $S=I+L/2$, where $I$ is the identity matrix and $L$ is the Laplacian matrix of the graph. The off-diagonal entries of $L$ are $L_{ij}=-1$ if the vertices $i$ and $j$ are connected by an edge and $L_{ij}=0$ otherwise. The diagonal entries $L_{ii}=d_{i}$ are the vertex degrees, the numbers of edges that meet at the vertices $i$.
@@
For a matrix $T$ that is assigned to an undirected graph with $m$ vertices as described above, the expected value and the variance (14) are
$E=1,\quad V=\frac{2g_{2}+6g_{1}}{(g_{1}^{2}+4g_{1}+4)\,m+(4g_{1}+8)},$ (26)
where $g_{1}$ and $g_{2}\geq g_{1}^{2}$ are the mean values of the vertex degrees and their squares:
$g_{1}=\frac{1}{m}\sum_{i=1}^{m}d_{i},\quad g_{2}=\frac{1}{m}\sum_{i=1}^{m}d_{i}^{2}.$ (27)
Proof
By (24), in the given case the constants (15) possess the representation
$A_{1}=\frac{g_{1}+2}{2}\,m,\quad A_{2}=\frac{g_{2}+5g_{1}+4}{4}\,m$
and it is $n-m=mg_{1}/2$. Therefore, the proposition follows from (16). ∎
For the matrices assigned to a family of graphs for which the above two mean values remain bounded independent of the number of vertices, the variance tends to zero as the number of vertices goes to infinity. The matrices assigned to graphs whose vertices up to one are connected with a designated central vertex, but not with each other, form the other extreme. For these matrices, the variances decrease to a limit value greater than zero as the number of vertices increases. However, such matrices are a rare exception, not only with respect to the above random matrices, but also in the context of matrices assigned to graphs. Because of
$V<\frac{2g_{2}}{g_{1}^{2}}\,\frac{1}{m},$ (28)
the inequality $V\geq\delta$ implies the lower bound $g_{2}/g_{1}^{2}>m\delta/2$. Consider the random graphs with a fixed number $m$ of vertices that are with a given probability $p$ connected by an edge, or those with a correspondingly given number of edges. Sampling the variances assigned to a large number of graphs in such a class, one sees that these variances do not exceed the value $2/m$ with a high probability.
With the exception of a few extreme cases, it can be observed that the distribution of the values $\|T^{t}\eta\|^{2}$ for points $\eta$ on the unit sphere of $\mathbb{R}^{n}$ more and more approaches the normal distribution with the expected value and the variance (14) as the dimensions increase, a fact that underlines the importance of these quantities. This can be verified by evaluating the expression $\|T^{t}\eta\|^{2}$ at a large number of independent and uniformly distributed points $\eta$ on the unit sphere and comparing the frequency distribution of the resulting values with the given Gauss function
$\frac{1}{\sqrt{2\pi V}}\,\exp\left(-\,\frac{(\,t-E)^{2}}{2V}\,\right).$ (29)
@@
for a large number of vectors $x\in\mathbb{R}^{m}$ and $y\in\mathbb{R}^{n-m}$ with independent and standard normally distributed components. The squares of the euclidean norms $\|y\|$ then follow the $\chi^{2}$-distribution with $n-m$ degrees of freedom. Therefore, their calculation can be replaced by the calculation of a single scalar quantity. The amount of work then remains proportional to the dimension $m$, no matter how much the dimensions differ and how large their difference $n-m$ is. Let $T$ be the matrix assigned to the graph associated with the $\mathrm{C}_{60}$-fullerene molecule, which consists of the ninety edges of a truncated icosahedron and its sixty corners as vertices. The degree of these vertices is three and the assigned variance therefore $V=9/380$. The frequency distribution of the values $\|T^{t}\eta\|^{2}$ for a million randomly chosen points $\eta$ on the unit sphere and the corresponding Gauss function (29) are shown in Fig. 2.